OpenStudy (anonymous):

It takes 1 pound of flour to make y cakes. The price of flour is w dollars for x pounds. In terms of w, x and y, what is the dollar cost of the flour required to make 1 cake?

OpenStudy (anonymous):

A. xy/w B. y/wx C. w/xy D. wx/y E. wxy

OpenStudy (solomonzelman):

Lets just simply write the data for convenience: \(\large\color{black}{ \displaystyle 1{~} {\rm pound}= y{~} {\rm cakes} }\) \(\large\color{black}{ \displaystyle {\rm $}w= x{~} {\rm pounds} }\) ------------------------------- 1 cake is ?

OpenStudy (anonymous):

make it up?

OpenStudy (anonymous):

5 cakes?

OpenStudy (solomonzelman):

excuse me, the cost for 1 cake (not flour)

OpenStudy (solomonzelman):

lets do it as it is: \(\large\color{black}{ \displaystyle {\rm $}w= x{~} {\rm pounds} }\) \(\large\color{black}{ \displaystyle {\rm $}\frac{w}{x}= 1{~} {\rm pounds} }\) you know that: `1 pound=y cakes` and thus, \(\large\color{black}{ \displaystyle {\rm $}\frac{w}{xy}= \frac{1}{y} {~} {\rm pounds}=1{~\rm cake} }\)

OpenStudy (solomonzelman):

because if `1 pounds = y cakes` then, `1/y pounds = 1 cake`